OpenStudy (anonymous):

Timmy writes the equation f(x) = 1/4x – 1. He then doubles both of the terms on the right side to create the equation g(x) = 1/2x – 2. How does the graph of g(x) compare to the graph of f(x)?

OpenStudy (anonymous):

@peachpi

OpenStudy (triciaal):

compare with y = m x + b m1 = 1/4 and m2 = 1/2 g(x) is steeper b1 = -1 and b2 = -2 = b1 - 1 graph moved down 1 unit

OpenStudy (triciaal):